z(x+y)p+z(x-y)q=x²+y² where p=dz/dx
q=dz/dx (partial differentiation) (i couldn't find its symbol so used d only)
1
Expert's answer
2020-12-28T13:27:34-0500

P=z(x+y)=zx+zy

Q=z(x-y)=zx-zy

R=x2+y2

"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

"\frac{dx}{zx+zy}=\frac{dy}{zx-zy}=\frac{dz}{x^2+y^2}"

Multiplyers: x,-y,-z

xdx-ydy-zdz=0

"\frac{x^2}{2}-\frac{y^2}{2}-\frac{z^2}{2}=C_1"

Multiplyers: -x,y,z

-xdx+ydy+zdz=0

"\frac{-x^2}{2}+\frac{y^2}{2}+\frac{z^2}{2}=C_2"